OpenStudy (anonymous):

find the derivative of f(x)=(cosx/1-sinx)^2 steps to finding this answer?

OpenStudy (anonymous):

Apply the chain rule followed by the quotient rule.

OpenStudy (anonymous):

Where are you getting stuck?

OpenStudy (anonymous):

On the chain rule part

OpenStudy (anonymous):

Okay. The chain rule is \((f\circ g)'(x)=f'(g(x))\cdot g'(x) \), right? So, we need to figure out what \(f\) and \(g\) are in this problem. If \(f(x)=\left(\dfrac{\cos x}{1-\sin x}\right)^2\), we can take \(f(x)=x^2\) and \(g(x)=\dfrac{\cos x}{1-\sin x}\). Do you see how that part works?

OpenStudy (anonymous):

If \(f(x)=x^2\) and \(g(x)=\dfrac{\cos x}{1-\sin x}\), then \(f(g(x))=\left(\dfrac{\cos x}{1-\sin x}\right)^2\), our original formula.

OpenStudy (anonymous):

@dilfalicious Are you with me so far?

OpenStudy (anonymous):

first simplify f(x) so: f(x)=cos^2x/(1-sinx)^2 now recall cos^2x=1-sin^2x=(1-sinx)(1+sinx) f(x)=(1-sinx)(1+sinx)/(1-sinx)(1-sinx)=(1+sinx)/(1-sinx) now apply only quotient rule so f'(x)=((1-sinx)D(1+sinx)-(1+sinx)D(1-sinx))/(1-sinx)^2 =(1-sinx)(cosx)-(1+sinx)(-cosx)/(1-sinx)^2 =(cosx-sinxcosx+cosx+sinxcosx)/(1-sinx)^2 =2cosx/(1-sinx)^2 <answer

OpenStudy (anonymous):

The way @anonymoustwo44 did it will work, but you should also learn how to use the chain rule, because you'll need to use it often.

OpenStudy (anonymous):

what would be the final step?

OpenStudy (anonymous):

Which step is the final one depends on which order you do things in. What do you have so far?

OpenStudy (anonymous):

I'm not sure where to go from your last step you posted

OpenStudy (anonymous):

Well, you know now what \(f(x)\) and \(g(x)\) are, so you apply the Chain Rule formula. Find \(f'(g(x))\cdot g'(x) \)

OpenStudy (anonymous):

so the derivative of f times g(x)?

OpenStudy (anonymous):

No, the derivative of f of g of x, times the derivative of g of x. So you take the derivative of f(x) then plug in g(x), and multiply that by the derivative of g(x)

OpenStudy (anonymous):

I have to go to work, I'll be back on in about 8 hrs or so.

OpenStudy (anonymous):

\[(\cos2x/1-\sin2)\times g \prime(x) ?\]

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

Remember f(x) is just x^2, so f'(x) is just 2x, then substitute g(x) in for x, so 2g(x). Then mutiply by g'(x) Good luck, I'll be back on later